# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ Abstract
The success of Artificial Intelligence (AI) in multiple disciplines and vertical domains in recent years has promoted the evolution of mobile networking and the future Internet toward an AI-integrated Internet-of-Things (IoT) era. Nevertheless, most AI techniques rely on data generated by physical devices (e.g., mobile devices and network nodes) or specific applications (e.g., fitness trackers and mobile gaming). To bypass this circumvent, Generative AI (GAI), a.k.a. AI-generated content (AIGC), has emerged as a powerful AI paradigm; thanks to its ability to efficiently learn complex data distributions and generate synthetic data to represent the original data in various forms. This impressive feature is projected to transform the management of mobile networking and diversify the current services and applications provided. On this basis, this work presents a concise tutorial on the role of GAIs in mobile and wireless networking. In particular, this survey first provides the fundamentals of GAI and representative GAI models, serving as an essential preliminary to the understanding of the applications of GAI in mobile and wireless networking. Then, this work provides a comprehensive review of state-of-the-art studies and GAI applications in network management, wireless security, semantic communication, and lessons learned from the open literature. Finally, this work summarizes the current research on GAI for mobile and wireless networking by outlining important challenges that need to be resolved to facilitate the development and applicability of GAI in this edge-cutting area.
@@ I Introduction
Generative Artificial Intelligence (GAI), a prominent branch of Artificial Intelligence (AI) responsible for AI-Generated Content (AIGC), has recently attracted significant attention from the scientific community to big tech giants and governments. The three most popular representations of GAI are ChatGPT, Bing, and Google Bard, which are Large Language Model (LLM)-based chatbots, that can conversationally interact with humans and generate new data based on the input and contextual data [1]. Different from conventional AI and Machine Learning (ML) systems that are only trained on a certain dataset to create mapping abstractions to make highly accurate decisions, GAI platforms do not just learn from the input dataset but also rely on the interaction of data to combine, modify, and create new information to reinforce its learning process. Driven by consumer interests, GAIs have recently emerged as a powerful tool for various uses. For example, in robotics control, Reinforcement Learning (RL) is a key technique for enabling robots to learn from their own experiences, which are then translated into appropriate movements and actions in real time. Through GAI interaction with RL, robots become more humanoid in generating new motions and movement skills without human intervention or initial knowledge provision [2]. Another example is in the field of cybersecurity for the Internet-of-Things (IoT), Deep Generative Models (DGM), fundamental components behind GAI, help generate patterns and address critical challenges involving the lack of malicious samples and the rapid evolution of cyberattacks [3]. As a consequence, the adoption of GAIs is projected to produce new ways to significantly improve current AI-enabled algorithms in mobile networking and the IoT.
Figure 1: Illustration of a high-level GAI architecture.
Fig. 1 illustrates a high-level GAI architecture, where the paradigm is composed of three main blocks. The “grey” block refers to the GAI models, in which the selection and optimization depend on use cases, relevant data, and performance indicators. Pretrained LLMs and Generative Adversarial Networks (GANs) are two representative GAI models that have found a lot of success in real GAI-enabled applications, such as chatbots, code generation, search engines, and cybersecurity. The “orange” block indicates the system settings used as the input to the GAI models to generate the output data. More precisely, users can provide initial guidance about the data to be generated and the links between the input data and expected output. For example, a typical user can input “what is generative AI?” and “what are essential blocks of GAI?” if they seek to receive draft answers from GAI tools. Finally, the “green” block corresponds to data generated from GAI models based on input from users (e.g., text, audio, video, and code), which then are used to provide other services, such as enhanced network management and improved wireless security in mobile communication and wireless networks. Fig. 2 shows a comprehensive overview of popular GAI developer tools and services, showcasing the layered structure from foundational hardware to essential cloud services and advanced model safety tools.
Figure 2: Popular developer tools and services for building GAI frameworks.
@@ I-A State-of-the-Arts
AI techniques have become widespread in various research areas in mobile networking and the IoT, such as network management, wireless security, resource allocation, and edge services. In recent years, data-driven and AI-enabled approaches have established clear performance superiority over many conventional optimization-based designs, which are limited primarily by network scalabilities in handling massive connections and heterogeneous variations [4]. Additionally, AI techniques enable efficient approximation of computationally challenging algorithms, to effectively overcome nonlinear and unknown patterns in wireless networks, and rapidly accelerate existing model-based algorithms. On this basis, there have been several surveys of different AI paradigms and their applications for mobile networking, such as Deep Learning (DL) [5, 6], Deep Reinforcement Learning (DRL) [7, 8], Federated Learning (FL) [9, 10], and swarm intelligence [11]. However, current AI-based approaches for mobile and wireless networking face several issues. Specifically, conventional AI algorithms usually require a large volume of labeled training data to facilitate better performance. Unfortunately, collecting and labeling training data in complex environments like wireless networks are costly, time-consuming, and require intensive human interventions. In addition, with an enormous number of diverse users, future wireless systems are highly dynamic and uncertain. It can significantly degrade the performance of conventional AI models, which are usually trained in a specific environment/condition. Although the distributed learning paradigm can help to facilitate the learning process of AI-based approaches, it still suffers heavily from scarcity of data required for efficient training [12]. Moreover, due to the differences in local system characteristics, the distributed AI system suffers from the long-tailed distributed data phenomenon. This data imbalance results in a negative transfer toward the data cluster with a dominant size. To address all these practical challenges, GAI is considered an efficient solution thanks to its properties of domain adaptation, data generation, and abnormal detection. Aiming to address these challenges, there are several surveys in the literature focusing on the applications of GAI in various domains such as mobile networking, Semantic Communication (SemCom), and security and privacy, as summarized in Table I.
@@
For example, GANs, a notable GAI technique, are reviewed in several recent works [13, 14, 15, 16, 17, 18], providing preliminary discussions about the use of this GAI-specific technique in emerging networks. More recently, GAI has been reviewed for several research issues in wireless networks. For instance, in [19], the authors provide a detailed survey of training solutions for GANs that deal with challenges of model collapsing and gradient vanishing, providing meaningful techniques and guidance for mobile network deployments. Similarly, in [20], the authors analyze GANs with various stability issues and survey several training solutions by assessing the obstacles. In [21], several critical GAN analyses of computer and communication networks are outlined, with a focus on network management, security, and mobile networks. Differently, the authors in [22] introduce the concept of on-device GAI, where multiple devices, functioning as distinct learning agents, deploy on-device GAI models for the collaborative processing of complex tasks. In [23], a direction on GAI services at the network edge has been outlined in various use cases, followed by guidance on building a hierarchical collaboration architecture of IoT devices, mobile edges, and centralized clouds. In [24], wireless network management efficiency can be improved by DL models in general and particular applications of DGM. In addition, the authors in [25] present a review of the applications of GAI for physical layer communications, focusing on emerging problems such as signal detection, physical layer security, and Joint Source-Channel Coding (JSCC). Considering a different perspective, the authors in [26] discuss the potential of common GAI techniques, e.g., GANs, VAEs, and normalizing flows, in addressing the design challenges of unmanned vehicle swarms.
@@
On another front, the interplay of GAI and SemCom has opened up a new avenue for reducing resource consumption while maintaining the reliability of wireless communication networks [27]. Fundamentally, SemCom relies on two fundamental techniques [28, 29] of knowledge extraction and lossy data compression, both rooted in AI. As such, the current landscape of SemCom designs represents a paradigm shift, with these innovative concepts emerging across various disciplines. In the literature, several surveys aim to elucidate the intricate connections between the potential AI techniques within GAI and their application in the field of SemCom. From the perspective of Knowledge Graph (KG), the current SemCom is based on the KG embeddings from Natural Language Processing (NLP), which are comprehensively reviewed in [30, 31, 32, 33]. From the perspective of lossy compression, despite a huge range of research that has both integrated and without knowledge representation to improve data reconstruction performance, no studies have surveyed these techniques in general and lossy compression-aided SemCom in particular.
@@
In addition to the aforementioned dominant features of GAI, it also possesses significant merits in the cutting-edge area of security. For instance, in [34], Zhang et al. explore GANs in transportation, including autonomous driving, data completion, and traffic anomaly detection. In [35], Liu et al. explore how adversarial ML can be both a threat and a defense mechanism in mobile networks. In [36], Cai et al. provides a discussion of the potential benefits and challenges of GANs in the context of privacy and security in mobile networks. In [37], Deldjoo et al. present the role and application of GANs in improving data distribution learning of recommendation systems. In [38], Dunmore et al. survey and discuss the potential of GAN applications in enhancing cybersecurity in mobile networks, particularly in Intrusion Detection Systems (IDS). In [39], Lin et al. investigate and inspect the use of GANs in enhancing trajectory prediction, security, and anomaly detection of the wireless-based transportation domain. However, a recent investigation in [40] points out that the use of GAI exposes potential security vulnerabilities, and this will generate a double-edged sword for wireless security if GAI is not exploited properly.
@@ I-B Motivations and Contributions
A common idea in existing studies is that most focus is on clarifying either applications or challenges. The increasing emphasis on generative AI technologies is expanding into various applications and domains within mobile and wireless networking. However, current studies do not provide comprehensive GAI evaluation and its applications. To the best of the authors’ knowledge, the fundamentals and applications of GAI as outlined above have not been thoroughly examined in the literature. Consequently, we aim to bridge this gap by providing the necessary preliminaries to GAI and its key applications in the future of the IoT and mobile networking. This work centralizes all recent developments in GAI and complementary technologies and remarkable advancements in various aspects of mobile networks. At the same time, this work also seeks to gather the recent works in this trending research field and provide a unified repository. The notable contributions of our work are listed below.
@@
Table I: Existing survey papers about GAI for mobile networking. ✓, ✗, and ✱  indicate that the topic is well-covered, uncovered, and partially covered, respectively.
Ref. Year Focus ML GAI NetMan Security SemCom
[14] 2022 Assistance of GAN in anomaly detection, with analysis techniques and application domains.
[15] 2022 Generative adversarial learning for secure, and reliable trust management in real-time communications.
[16] 2023 AIGC for efficient and accessible content creation using large-scale models with unimodal and multimodal interactions.
[17] 2023 Focus on AIGC model architecture to output types, in tasks like text, images, and videos.
[19] 2021 Surveying training solutions for GAN with mode collapse, and gradient vanishing.
[20] 2021 Analyzes GANs with various stability issues, and training solutions by assessing obstacles.
[21] 2021 Explores aspects of computer and communication networks, including network management, security, and mobile networks.
[22] 2023 Integration of multi-agent GAI in wireless networks based on cloud-based LLMs.
[23] 2024 Personalized AIGC services at mobile edge networks ensuring privacy measures.
[24] 2023 Applications of deep generation models in improving the efficiency of wireless network management.
[27] 2023 Architecture, challenges, and outlook of Generative AIs.
[34] 2019 Explores GAN in transportation including autonomous driving, data completion, and traffic anomaly detection.
[36] 2021 Focuses on GAN for privacy and security contexts, highlighting potential benefits and challenges.
[37] 2021 Security of recommender systems and its application in GANs for data distribution learning.
[39] 2023 Trajectory prediction, security, and anomaly detection in the transportation domain.
[42] 2023 Causal SemCom for digital twin-based wireless systems.
[43] 2023 FL solutions to enhance the efficiency and privacy protection of AIGC using stable diffusion models.
[44] 2023 AaaS architecture that deploys AIGC models in edge networks.
This Work - A review on applications of GAI for network management, security, and SemCom in mobile networking.
I-C Paper Organization
In the study, Section II presents a preliminary to GAI, including the development history, data-processing and data-creation techniques, and representative models. In Section III, we comprehensively review the role of GAI in mobile network management. Section IV reviews the use of GAI in various wireless security issues, including data obscurity, intrusion detection, jamming attacks, and generative steganography, while Section V describes the applications of GAI for research advances in SemCom. Section VI discusses crucial challenges and explores potential research directions to facilitate applications of GAI techniques. We finally conclude the paper in Section VII.
II Generative AI: Preliminary
This section introduces a preliminary to GAI, along with its structure and data-creation techniques. It lays the platform for building robust applications of GAI in mobile networking and the future Internet.

### v2

@@ Abstract
The success of Artificial Intelligence (AI) in multiple disciplines and vertical domains in recent years has promoted the evolution of mobile networking and the future Internet toward an AI-integrated Internet-of-Things (IoT) era. Nevertheless, most AI techniques rely on data generated by physical devices (e.g., mobile devices and network nodes) or specific applications (e.g., fitness trackers and mobile gaming). Therefore, Generative AI (GAI), a.k.a. AI-generated content (AIGC), has emerged as a powerful AI paradigm; thanks to its ability to efficiently learn complex data distributions and generate synthetic data to represent the original data in various forms. This impressive feature is projected to transform the management of mobile networking and diversify the current services and applications provided. On this basis, this work presents a concise tutorial on the role of GAIs in mobile and wireless networking. In particular, this survey first provides the fundamentals of GAI and representative GAI models, serving as an essential preliminary to the understanding of GAI’s applications in mobile and wireless networking. Then, this work provides a comprehensive review of state-of-the-art studies and GAI applications in network management, wireless security, semantic communication, and lessons learned from the open literature. Finally, this work summarizes the current research on GAI for mobile and wireless networking by outlining important challenges that need to be resolved to facilitate the development and applicability of GAI in this edge-cutting area.
@@ I Introduction
Generative Artificial Intelligence (GAI), a prominent branch of Artificial Intelligence (AI) responsible for AI-Generated Content (AIGC), has recently attracted significant attention from the scientific community to big tech giants and governments. The three most popular representations of GAI are ChatGPT, Bing, and Google Bard, which are Large Language Model (LLM)-based chatbots that can conversationally interact with humans and generate new data based on the input and contextual data [1]. Different from conventional AI and Machine Learning (ML) systems that are only trained on a certain dataset to create mapping abstractions to make highly accurate decisions, GAI platforms do not just learn from the input dataset but also rely on the interaction of data to combine, modify, and create new information to reinforce its learning process. Driven by consumer interests, GAIs have recently emerged as a powerful tool for various uses. For example, in robotics control, Reinforcement Learning (RL) is a key technique for enabling robots to learn from their own experiences, which are then translated into appropriate movements and actions in real time. Through GAI interaction with RL, robots become more humanoid in generating new motions and movement skills without human intervention or initial knowledge provision [2]. Another example is in the field of cybersecurity for the Internet-of-Things (IoT), Deep Generative Models (DGM), fundamental components behind GAI, help generate patterns and address critical challenges involving the lack of malicious samples and the rapid evolution of cyberattacks [3]. Consequently, adopting GAIs is projected to produce new ways to significantly improve current AI-enabled algorithms in mobile networking and the IoT.
Figure 1: Illustration of GAI in the context of wireless IoT networks.
Fig. 1 illustrates GAI in the context of IoT, where its architecture consists of six layers, each dedicated to specific functions [4]. The foundation layer, Infrastructure, encompasses computing resources based on GPUs, CPUs, or distributed systems, data storage solutions such as data lakes, databases, data warehouses, and network connections such as high-speed networks and cloud services. Architecture, a critical aspect of GAI’s structural design, includes the optimal selection and configuration of neural networks for specific generative tasks. It also assesses the models’ capability and efficiency in learning from data and generating new content, utilizing key components like Transformers and Generative Adversarial Networks (GANs). Herein, Transformers use Attention mechanisms to manage dependencies between input and output data, while GANs generate high-quality synthetic data. The third layer, Machine Learning (ML) foundation model, involves building ML models trained on vast datasets to capture fundamental patterns, including training data, algorithms, and model evaluation, making GAI powerful in learning and generalizing from data. The LLM layer focuses on implementing LLMs using model architectures (e.g., Transformer, GPT, BERT, T5), training data with considerations of text corpora, diversity, and representation, and fine-tuning for task-specific adaptation or transfer learning. The Application Programming Interface (API) layer provides standardized interfaces for accessing and utilizing GAI models, including Representational State Transfer Application APIs (RESTful APIs) for Hypertext Transfer Protocol (HTTP) activities, Graph Query Language (GraphQL) APIs for client requests and interactions, and software development kits (SDKs) for programming. Finally, the Application layer bridges GAIs to real-world scenarios across various domains, showcasing the tangible benefits of GAI in improving efficiency and driving innovation, such as in chatbots, code generation, search engines, and cybersecurity. More specifically, when a typical user inputs queries like ‘What is generative AI?” or “What are essential blocks of GAI?” to seek draft answers from GAI tools, the ‘grey’ block, representing trained GAI models, combines with the ‘orange’ block, which includes the system settings used as input to the GAI models, to generate output data in the form of text, audio, video, or code. The ‘green’ block then processes the data generated by the GAI models to respond to the user’s input. Thus, introducing GAI into not only mobile and wireless networking but also large-scale IoT will open exciting new research directions, paying the base to shift current network deployments close to the era of generative IoT [5].
@@ I-A State-of-the-Arts
AI techniques have become widespread in various research areas in mobile networking and the IoT, such as network management, wireless security, resource allocation, and edge services. In recent years, data-driven and AI-enabled approaches have established superior performance over conventional optimization-based designs, which are limited primarily by network scalabilities in handling massive connections and heterogeneous variations [6]. Additionally, AI techniques enable efficient approximation of computationally challenging algorithms, to effectively overcome nonlinear and unknown patterns in wireless networks, and rapidly accelerate existing model-based algorithms. On this basis, there have been several surveys of different AI paradigms and their applications for mobile networking, such as Deep Learning (DL) [7, 8], Deep Reinforcement Learning (DRL) [9, 10], Federated Learning (FL) [11, 12], and swarm intelligence [13]. However, current AI-based approaches for mobile and wireless networking face several issues. Specifically, conventional AI algorithms usually require a large volume of labeled training data to facilitate better performance. Unfortunately, collecting and labeling training data in complex environments like wireless networks are costly, time-consuming, and require intensive human interventions. In addition, with an enormous number of diverse users, future wireless systems are highly dynamic and uncertain. It can significantly degrade the performance of conventional AI models, which are usually trained in a specific environment/condition. Although the distributed learning paradigm can help to facilitate the learning process of AI-based approaches, it still suffers heavily from scarcity of data required for efficient training [14]. Moreover, due to the differences in local system characteristics, the distributed AI system suffers from the long-tailed distributed data phenomenon. This data imbalance results in a negative transfer toward the data cluster with a dominant size. To address all these practical challenges, GAI is considered an efficient solution thanks to its properties of domain adaptation, data generation, and abnormal detection. Consequently, several surveys in the literature focus on the applications of GAI in various domains such as Network Management (NetMan), Semantic Communication (SemCom), and security and privacy, as summarized in Table I.
Table I: Existing survey papers about GAI for mobile networking.
 ●, ◐  and ○, indicate that the topic is well-covered, partially covered, and uncovered, respectively.
Ref. Year Focus ML GAI NetMan Security SemCom
[16] 2022 Analysis techniques and applications of GAN in anomaly detection.
[17] 2022 GAN for security and reliable trust management in real-time communications.
[18] 2023 AIGC for content creation using unimodal and multimodal large-scale models.
[19] 2023 AIGC models for output types in specific tasks (text, images, and videos).
[21] 2021 Training solutions for GAN with mode collapse, and gradient vanishing.
[22] 2021 Analyses of GANs’ stability, and training solutions by assessing obstacles.
[23] 2021 Computer and communication interplay (control, security, and mobile services).
[24] 2023 Integration of multi-agent GAI in wireless networks based on cloud LLMs.
[25] 2024 Personalized AIGC services for mobile edge networks’ privacy measures.
[26] 2023 Improves the efficiency of wireless network management by DGMs.
[27] 2024 GAI use cases across various industry mobile networks
[28] 2024 Data scarcity issues in 6G Wireless through GAI
[29] 2024 Enhanced physical-layer security of communication networks using GAI
[30] 2023 Architecture, challenges, and outlook of GAIs.
[31] 2019 GAN for autonomous driving, data completion, and traffic anomaly detection.
[33] 2021 GAN for privacy and security contexts, its benefits and challenges.
[34] 2021 Recommender systems’ security & GANs’ applications for distributed learning.
[36] 2023 Trajectory prediction, security, and anomaly detection in transportation sectors.
[37] 2024 GAN to enhance wireless network management
[39] 2023 Causal SemCom for digital twin-based wireless systems.
[40] 2023 Enhanced privacy of AIGC using stable diffusion models and FL solutions.
[41] 2023 AaaS architecture that deploys AIGC models in edge networks.
This Work - A review on applications of GAI for network management, security, and SemCom in mobile and wireless networking.
@@
For example, GANs, a notable GAI technique, are reviewed in several recent works [15, 16, 17, 18, 19, 20], providing preliminary discussions about the use of this GAI-specific technique in emerging networks. More recently, GAI has been reviewed for several research issues in wireless networks. For instance, in [21], the authors provide a detailed survey of training solutions for GANs that deal with challenges of model collapsing and gradient vanishing, providing meaningful techniques and guidance for mobile network deployments. Similarly, in [22], the authors analyze GANs with various stability issues and survey several training solutions by assessing the obstacles. In [23], several critical GAN analyses of computer and communication networks are outlined, with a focus on network management, security, and mobile networks. Differently, the authors in [24] introduce the concept of on-device GAI, where multiple devices, functioning as distinct learning agents, deploy on-device GAI models for the collaborative processing of complex tasks. In [25], a direction on GAI services at the network edge has been outlined in various use cases, followed by guidance on building a hierarchical collaboration architecture of IoT devices, mobile edges, and centralized clouds.
In [26], wireless network management efficiency can be improved by DL models in general and particular applications of DGM. In addition, the authors in [42] present a review of the applications of GAI for physical layer communications, focusing on emerging problems such as signal detection, physical layer security, and Joint Source-Channel Coding (JSCC). Considering a different perspective, the authors in [43] discuss the potential of common GAI techniques, e.g., GANs, Variational Autoencoders (VAEs), and normalizing flows, in addressing the design challenges of unmanned vehicle swarms. In [27], the authors overview the classification of GAI solutions used in mobile networks and outline the involved performance requirements. The authors in [28] discuss the use cases of discriminative AI in Sixth-Generation (6G) applications and services. They then detail how GAI enhances or complements discriminative AI models across various domains, including physical layer design, network management, cross-layer network security, and localization and positioning. Based on these analyses, they conclude by emphasizing GAI’s core roles in cutting-edge areas such as massive Multiple-Input Multiple-Output (MIMO), THz communications, integrated sensing and communications, near-field communications, digital twins, SemCom, AI-generated content services, mobile edge computing and edge-AI, adversarial ML, and trustworthy AI. In [29], the authors provide a tutorial and case study of using deep generation models for wireless network management, followed by guidelines on how GAI overcomes major issues of training data scarcity, limited flexibility, and dynamic network states.
@@
On another front, the interplay of GAI and SemCom has opened up a new avenue for reducing resource consumption while maintaining the reliability of wireless communication networks [30]. SemCom relies on two fundamental techniques [44, 45] of knowledge extraction and lossy data compression, both rooted in AI. The current landscape of SemCom designs represents a paradigm shift, with these innovative concepts emerging across various disciplines. In the literature, several surveys aim to elucidate the intricate connections between the potential AI techniques within GAI and their application in the field of SemCom. From the perspective of Knowledge Graph (KG), the current SemCom is based on the KG embeddings from Natural Language Processing (NLP), which are comprehensively reviewed in [46, 47, 48, 49]. From the perspective of lossy compression, no studies have surveyed these techniques in general and lossy compression-aided SemCom in particular, despite a huge range of research that has both integrated and without knowledge representation to improve data reconstruction performance.
@@
In addition to the aforementioned dominant features of GAI, it also possesses significant merits in the cutting-edge area of security. For instance, in [31], Zhang et al. explore GANs in transportation, including autonomous driving, data completion, and traffic anomaly detection. In [32], Liu et al. explore how adversarial ML can be both a threat and a defense mechanism in mobile networks. In [33], Cai et al. provide a discussion of the potential benefits and challenges of GANs in the context of privacy and security in mobile networks. In [34], Deldjoo et al. present the role and application of GANs in improving data distribution learning of recommendation systems. In [35], Dunmore et al. survey and discuss the potential of GAN applications in enhancing cybersecurity in mobile networks, particularly in Intrusion Detection Systems (IDS). In [36], Lin et al. investigate and inspect the use of GANs in enhancing trajectory prediction, security, and anomaly detection of the wireless-based transportation domain. Similarly, Zhao et al. in [37] also emphasize the role of advanced GAI approaches in dynamically enhancing the security, authentication, availability, resilience, and integrity of communications in the physical layer of communication networks where traditional AI approaches often struggle with the evolving physical properties of the transmission channels and emerging cyber threats. However, a recent study in [50] points out that the use of GAI exposes potential security vulnerabilities, and this will generate a double-edged sword for wireless security if GAI is not exploited properly.
@@ I-B Motivations and Contributions
It can be observed that existing surveys mostly focus on specific applications of GAI on the challenges of integrating GAI into mobile and wireless networking. However, current studies do not provide comprehensive evaluation, roles, and applications of GAI in mobile and wireless networking (i.e., GAI for networks). To the best of the authors’ knowledge, the fundamentals and applications of GAI as outlined above have not been thoroughly examined in the literature. Consequently, we aim to bridge this gap by providing the necessary preliminaries to GAI and its key applications in the future of IoT and mobile networking. This work centralizes all recent developments in GAI and complementary technologies and remarkable advancements in various aspects of mobile networks. Moreover, this work also seeks to gather the recent trending research in the field and provide a unified repository. The notable contributions of our work are listed below.
@@
Figure 2: The outline of this survey paper.
I-C Structure of the Survey
The structure of this survey is summarized in Fig. 2. Specifically, we start with a preliminary of GAIs in Section II, covering the development history, data-processing and data-creation techniques, and representative models. Then, we delve into the role details of GAI in mobile network management (Section III), with emphasis on levels of network control, resource allocation, network routing, and channel estimations. Next, we discuss the need for GAIs to enhance wireless security networks (Section IV) in terms of data obscurity, intrusion detection, jamming attacks, and generative steganography. Moreover, we present the promising applications of GAIs in a new communication era of mobile networks and SemCom in Section V. Finally, we summarize the challenges and potential research directions of integrating GAIs into mobile networking in Section VI, and we conclude the paper in Section VII.
II Preliminary of Generative AI
This section presents a preliminary of AI, including its evolutionary history, primitive structure, multi-modal functionality, and representative models. It serves as a basic guide for general readers to understand the following sections of GAI applications. For ease of understanding, Fig. 3 summarizes all information related to GAI.